The following graph plots the current security market line (SML) and indicates the return that investors require from holding stock from Happy Corp. (HC). Based on the graph, complete the table that follows. REQUIRED RATE OF RETURN (Percent] 20.0 16.0 12.0 Return on HCs Stock 4.0 0.5 1.5 2.0 RISK (Beta) 0.0 1.0 CAPM Elements Value Risk-free rate (rRF) Market risk premium (RPM) Happy Corp. stocks beta Required rate of return on Happy Corp. stockAn analyst believes that inflation is going to increase by 2.0% over the next year, while the market risk premium will be unchanged. The analyst uses the Capital Asset Pricing Model (CAPM). The following graph plots the current SML. Calculate Happy Corp. s new required return. Then, on the graph, use the green points (rectangle symbols) to plot the new SML suggested by this analysts prediction. Happy Corp.s new required rate of return is Tool tip: Mouse over the points on the graph to see their coordinates. REQUIRED RATE OF RETURN (Percent 20 New SML 16 12 0.0 0.40.8 12 .6 2.0 RISK (Beta) Clear AllThe SML helps determine the risk-aversion level among investors. The flatter the slope of the SML, the the level of risk aversion. higher lower Which of the following statements best describes a shift in the SML caused by increased risk aversion? O The risk-free rate will increase. The risk-free rate will remain constant. The risk-free rate will decrease.
Answer #1

Based on the graph :

The risk free return is = 4%

From the graph, when the beta =1 , the market return is about 10%. So, Market risk premium = market return - risk free return = 10-4 = 6%

Market risk premium = 6%

Happy Corp Stock Beta = 0.6

Required return on happy corp. stock = 4 + 0.6*(10-4) = 7.6%

Next year when inflation increases by 2%, risk free return = 6% and Market risk premium = unchanged = 6%

So, Happy corp's new required rate of return = 6 + 0.6*6 = 9.6%

The new SML in green is as shown below:

Flatter the slope of the SML, the lower the level of risk aversion

which of the following best describes a shift in the SML caused by increased risk aversion?

Risk free rate will increase (Option A)